Brianne is three times as old as Camila. Four years from now she will be just two times as old as Camila. How old are Brianne and Camila now?
Answer by ankor@dixie-net.com(22740)

b = B's present age
c = C's present age
:
Write an equation for each statement:
:
"Brianne is three times as old as Camila."
b = 3c
:
" Four years from now she will be just two times as old as Camila."
b + 4 = 2(c+4)
b + 4 = 2c + 8
b = 2c + 8 - 4
b = 2c + 4
Replace b with 3c
3c = 2c + 4
3c - 2c = 4
c = 4 yrs is C's present age
:
I know you can find B's age now, check it in the 4 yr equation